A truck driver named gabriel went crazy and decided to ride off the edge of a cliff. Just at the edge, his velocity is horizontal with magnitude of 12.0m/s. Find his position, distance fromthe edge of the cliff. Find the velocity after .50 seconds.
1
Expert's answer
2019-08-08T10:56:24-0400

t = 0.50 s

"v_{0}=12 m\/s"

This is a question on projectile motion. 

The distance from the cliff's edge after truck leaves is the range. 

"range=v_{0}t"

Where "v_{0}" = initial velocity

t = time

Range = 0.5 × 12 = 6 m

For the vertical components :

We will use the following equation of motion :

"v=v_{0}+at"

a = acceleration due to gravity

"v" = final velocity

"v=12+5=17 m\/s"